A 60 mililiter solution has 50% acid. How much pure water is needed to get 30% acid?
Answer by checkley71(8403)

.5*60=.3(60+x)
30=18+.3x
.3x=30-18
.3x=12
x=12/.3
x=40 ml. of water is needed.
PROOF:
.5*60=.3(60+40)
30=.3*100
30=30